OpenStudy (anonymous):

How are the functions y = x and y = x – 3 related? How are their graphs related?

OpenStudy (anonymous):

A.Each output for y = x – 3 is 3 more than the corresponding output for y = x. The graph of y = x – 3 is the graph of y = x translated down 3 units. B.Each output for y = x – 3 is 3 more than the corresponding output for y = x. The graph of y = x – 3 is the graph of y = x translated up 3 units. C.Each output for y = x – 3 is 3 less than the corresponding output for y = x. The graph of y = x – 3 is the graph of y = x translated down 3 units. D.Each output for y = x – 3 is 3 less than the corresponding output for y = x. The graph of y = x – 3 is the graph of y = x translated up 3 units.

OpenStudy (dape):

They are parallell, the second line is always 3 units below the first line. Here is a graph: 

OpenStudy (anonymous):

but i still dont get it:/

OpenStudy (anonymous):

try out each possible answer and see if they turn out to be true

OpenStudy (dape):

That two objects are translated in space just means that they are moved. So that the line is translated down 3 units is the same as moving the first line down 3 steps. If you think about it you should be able to figure it out.

OpenStudy (anonymous):

y = x - 3 is obtained when you shift y = x down 3 units !